OpenStudy (anonymous):

Which point lies on a circle with a radius of 5 units and center at P(6, 1)?

OpenStudy (anonymous):

There are an infinite number of points on the circle. So, find the equation of the circle and you will be able to find as many points as you wish. Keep in mind that the equation of the circle would be (x - h)^2 + (y - k)^2 = r^2 where r = radius of the circle and (h,k) is the center of the circle.

OpenStudy (anonymous):

answers to choose From A:) Q(1, 11) B:)R(2, 4) C;) S(4, -4) D;)T(9, -2)

OpenStudy (anonymous):

If you are given choice, then complete the problem by finding the distance from the center of the circle to each choice, using the distance formula, and see where you get a distance of 5.

OpenStudy (anonymous):

the closest i get to is B

OpenStudy (anonymous):

What is the distance from the center of the circle to (2,4)? What did you get using the Distance Formula?

OpenStudy (anonymous):

You should get an EXACT distance of 5 using choice (B), not close, as you said.

OpenStudy (anonymous):

can youtell me if its B please

OpenStudy (anonymous):

It's B